OpenStudy (anonymous):

The force of gravity, F, exerted between two objects is equal to the product of the gravitational constant, G, the mass of the first object, m1, and the mass of the second object, m2, divided by the square of the distance between their centers, d. This is often used to determine the gravitational attraction between two massive bodies, such as planets, in space. a. Write the formula for F described in the situation above.

OpenStudy (johnweldon1993):

So the force of gravity "F" is the product of "G" "m1" and "m2" all divided by the square of the distance between centers "d" So it looks like \[\large F = \frac{G\times m_1 \times m_2}{d^2}\]